OpenStudy (anonymous):

What is the greatest common factor (GCF) of 84 and 96? A. 12 B. 32 C. 21 D. 6

OpenStudy (anonymous):

A. 12 :)

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

To find the GCF, you take only common factors with the smaller exponent. Then multiply them together. From 2^2 and 2^5, use 2^2. From 3 and 3, use 3. From 7, do not take any since it's not a common factor. GCF = 2^2 * 3 = 4 * 3 = 12